OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 3x2 − y3 − y3 − z if x = 3, y = −2, and z = −5.

OpenStudy (nottim):

What do you need help with.

OpenStudy (anonymous):

to find the answer

OpenStudy (nottim):

Lol. Just sub in the values.

OpenStudy (nottim):

you can easily do that with a claculator. and bedmas/pedmas

OpenStudy (anonymous):

= 3*3^2 − 2*(−2)^3 − 5 = 3*9 − 2*(−8) − 5 = 27 + 16 − 5 = 38

OpenStudy (anonymous):

or Take your calculator. Everywhere you see and x, replace it with (3) 3 in brackets. Same for y and z. and also By hand. Work out x^2, y^3. and replace them in the equation with your result (in brackets). Remember that 27 - (-8) will change to 27 + 8 etc.

OpenStudy (jack1):

all good now logan?